# GraphAr: An Efficient Storage Scheme for Graph Data in Data Lakes

### 18 Dec 2023 v2

@@ 1 Introduction
In sync with these trends, graph data have gained increasing importance, especially for modeling complex relationships. Leading graph databases like Neo4j [13], TigerGraph [31], and JanusGraph [10] leverage the Labeled Property Graph (LPG) model [20, 23, 21] for this purpose. The recent ISO SQL:2023 standard includes a SQL/PGQ extension that not only facilitates querying property graphs but also enables the creation of property graph views from relational tables [19]. This groundbreaking inclusion highlights the growing convergence of relational and graph data models and emphasizes the need for integrating LPGs into data lakes. Consequently, LPGs are making their way into data lakes for multiple uses, from backups and archival storage for existing graph databases to natural extensions of transactional, log, and tabular data.
Managing and analyzing LPGs in data lakes offers several significant benefits. Graph-specific queries are often more naturally articulated in languages like Cypher [35], Gremlin [55], GQL [34], or SQL/PGQ, providing an intuitive framework for conducting comprehensive analysis of entity relationships, facilitating the discovery of valuable insights. Data lakes also offer the computational flexibility required for running complex graph algorithms, enabling the exploration of intricate patterns. Additionally, they provide cost-effective storage solutions, allowing organizations to utilize more affordable and colder storage options without sacrificing query performance. Most notably, data lakes enable seamless querying across both graph and relational data, ushering in a holistic approach to data analytics.
@@ 2.1 Tabular Formats in Data Lakes
Tabular data is key to data lakes, aiding efficient organization, analysis, and data extraction from large sets. Columnar formats like Parquet [5] and ORC [4] are popular due to their robust features. Unlike row-based formats such as CSV, they allow faster queries by enabling selective column reading, avoiding unnecessary data. Additionally, they offer diverse and efficient compression and encoding strategies, such as delta encoding to compress the variance between consecutive values, and run-length encoding to compress repetitive values. These techniques not only reduce storage needs but also enhance processing speeds.
Another advantage of Parquet and ORC is predicate pushdown, which enhances query performance by moving filters closer to the storage layer, thereby reducing subsequent data reads. The synergy of these attributes—selective column reading, efficient compression, and predicate pushdown—positions Parquet and ORC as the go-to choices for managing tabular data in data lakes. Previous studies have demonstrated the importance of leveraging their capabilities for optimizing relational data management [33, 40, 22]. Recent research [45, 60, 61, 62, 63, 6] has also explored enhancements to columnar formats, utilizing CPU instruction sets like BMI and SIMD. Without loss of generality, we will focus on Parquet in this paper, while the techniques discussed can be seamlessly adapted to other columnar formats such as ORC.
@@ 2.2 Labeled Property Graphs
Labeled Property Graphs (LPGs) [20, 23, 21] excel at representing complex relationships and semantics in a natural manner. Their flexible schema allows for accommodating the diverse and evolving nature of big data within repositories, making them ntegral to data lakes. LPG serves as the canonical data model in many graph systems [13, 36, 31] and graph query languages [59, 7, 35, 55, 34, 30], enabling queries and analytics to uncover valuable insights and patterns. Formally, an LPG is defined as $G=(V,E,T_{V},T_{E},P,L)$, where $V$ is a set of vertices, $E$ a set of edges, $T_{V}$ and $T_{E}$ the types of vertices and edges respectively, $P$ the properties, and $L$ the labels.
@@
Furthermore, each label $l\in L$ has a unique identifier, usually a string. Each vertex type $t_{v}$ is linked with a set of candidate labels $L(t_{v})\subseteq L$, allowing each vertex of this type to be assigned zero or more labels. Labels hold significant importance in LPGs as they represent classifications and characteristics of entities, whereas properties serve as attributes to store additional information. Although edge types could technically also be labeled, in this paper, we focus solely on vertex labels, aligning with common graph query practice  [^1] Graph query languages such as Cypher and Gremlin typically adhere to the convention that an edge can have one and only one classification, which corresponds to the edge type in the LPG model. Nevertheless, the strategies for vertex labels discussed in this paper can be seamlessly extended to accommodate edge labels..
@@ 2.3 Querying Labeled Property Graphs
These query-side inefficiencies highlight the limitations of using tabular formats for graphs. While they are proficient in storing and querying tabular data, they are inadequate for the specialized requirements of graph querying. The reliance on suboptimal workarounds, such as multiple joins for topology expansion and makeshift encoding schemes for label filtering, compromises query performance and complicates the query process. This paves the way for the challenges we seek to address in the following sections.
@@ 2.4 Key Challenges Addressed by GraphAr
The development of GraphAr is motivated by the specific limitations of existing tabular formats for both representing LPGs and supporting efficient graph queries. Below, we outline the key challenges that GraphAr aims to tackle, each corresponding to a dedicated section later in this paper.
@@
Each of these challenges represents a gap in the capabilities of current tabular formats for graph data, and they serve as the focus areas for the technical contributions of GraphAr detailed in the subsequent sections.
@@ 3.1 Goals and Non-Goals
Non-Goals. GraphAr does not intend to replace existing data lake formats like Parquet and ORC. Instead, but to maximize their benefits and offer additional features for LPGs.
@@ 3.2 Data Organization and Layout
Property columns, like pid and age, are named after their respective properties and hold the corresponding values with specified datatypes. In terms of labels, a set of candidate labels is defined for each vertex type. Then a vertex can have an arbitrary number of labels from the corresponding candidate set. For example, the vertex type Person may have labels to represent ethnicity and location. For efficient storage and filtering of labels, GraphAr uses a binary representation to maintain each label in an individual column named with angle brackets, e.g., <Asian> and <Enrollee>. Additionally, advanced encoding/decoding techniques are applied, which will be further discussed in Section 5.
Edge table. Edges are also organized and stored in Parquet files, similar to vertices. Figure 4 (c) showcases the layout of the edge table for type Person-Diagnosed-Disease, where Person and Disease represent the source and destination vertex types, while Diagnosed signifies the classification of the relationships. Each edge is associated with the internal IDs of its source and destination vertices, stored in columns named <src> and <dst>. Properties are stored in columns named according to their property names. Partitioning is also allowed in edges in a similar way to vertices.
@@ 3.3 Incorporation with Data Lakes
The design of GraphAr makes it especially well-suited for integration with data lakes, largely due to its reliance on widely adopted standards like Parquet for data storage and YAML for metadata management.
Data transformation and construction. The GraphAr format is essentially a specialized layout of Parquet files accompanied by a YAML metadata file. This enables the use of existing data processing frameworks like Apache Spark, Acero, and Hadoop, which can access various graph systems like Neo4j, TigerGraph and Nebula, or other types of database systems through their respective connectors. These frameworks can also ingest a multitude of data formats including logs, relational tables, JSON, and more. Such flexibility provides users with the ability to construct, transform, and store Labeled Property Graphs (LPGs) in data lakes from a wide array of data sources. To simplify the process of generating files in GraphAr, we have also provided a Spark library specifically designed for this purpose.
@@ 4.1 Workflow and Challenges
Given a column in vertex table that includes $m$ data pages, the page-aligned collections $\mathcal{C}=[C_{0},\ldots,C_{m-1}]$ is a list of up to $m$ collections. Each $C_{i}$ stores a set of internal IDs in the corresponding data page. Non-empty collections in $\mathcal{C}$ are retained, while empty ones are omitted.
@@ 4.2 Delta Encoding
Implementations based on Parquet. We utilize Parquet’s built-in support for delta encoding [43]. This technique is implemented based on miniblocks, with a default size of $32$ values. Each miniblock is binary packed using its own bit width, and for data alignment purposes, the bit width should be a power of $2$. The use of this delta encoding design allows us to adapt to changes in the data distribution, as the bit width of each miniblock is dynamically adjusted to minimize storage consumption. According to our evaluation across various real-world graphs, as detailed in Section 6.2, the delta encoding technique can reduce the expected loaded data volume (which can be measured by storage consumption) by $58.1\%$ to $81.0\%$ compared to without delta encoding. As a result, it brings an individual speedup of up to 3.6×3.6$\times$3.6 × for neighbor retrieval.
@@ 4.3 BMI-based Decoding
Another unsolved issue is the choice of implementing collection $C$ in PAC. A pioneering solution proposed in [45] underscores the transformation of indices into a bitmap representation to enable selection pushdown in columnar storages. To adopt this approach, to efficiently retrieve the properties of neighbors, which constitutes the second task of neighbor retrieval, we adopt a bitmap representation for each non-empty collection $C$ in PAC, while omitting empty collections. Specifically, we denote $B$ as a bitmap representation of a collection $C$, where $B[i]=1$ indicates the existence of $i$-th element. Figure 5 provides an example PAC and the bitmap representation, where only $C_{1}$ is non-empty, and $B_{1}$ is the bitmap representation of $C_{1}$. This bitmap representation can be used to facilitate the selection pushdown of vertex properties and labels, e.g., fetching the properties of name within the target vertex table Disease, for the neighbors of a person vertex.
@@
Intuition. For the sake of clarity, we initially consider a simplified scenario where each delta value is compressed to a 4-bit size. Conventionally, a two-step approach is used to decode the delta-encoded data, in which the current encoded value is added to the previously decoded ID to obtain the current ID, and then the bitmap is updated bit by bit based on the decoded IDs. However, our analysis reveals that this two-step process is redundant. By leveraging the bit-shifting encoding $1<<(d-1)$ for each delta value $d$, we can generate the bitmap by concatenating the bit-shifting encodings: $1<<(d_{n-1}-1)||\dots||1<<(d_{1}-1)||1<<(d_{0}-1)$, where $d_{0},d_{1},\dots,d_{n-1}$ represent $n$ delta values, and ||||| | represents the concatenation operator. This principle is visually depicted in Figure 6 (a).
@@
The subsequent challenge lies in generating the required mask. This can be accomplished with the $i$-th mask $m_{i}$ derived from the $i$-th bit-shifting encoding $s_{i}$ using the equation $m_{i}=(s_{i}<<1)-1$. The beauty of this operation is its potential for parallel execution, facilitated by direct manipulation of the mask sequence residing within a 64-bit register. Specifically, $M$ can be acquired via the following steps: 1) a bitwise (in our example, 16 bits) shift of each $s_{i}$ to the left by 1 bit, parallelized through SIMD instructions like _mm_slli_epi16; and 2) a bitwise subtraction of 1 from the result of the previous step, which can be accelerated via instructions like _mm_sub_epi16. All these SIMD instructions utilized in our implementations are widely available in modern CPUs, included in SSE2 (which we use) and more recent sets such as AVX2 and AVX-512.
@@ 5 Optimized Label Filtering
Labels serve as a representation of the classification or characteristics of vertices in a graph. Filtering vertices by labels is a fundamental syntax in graph query languages, as it allows querying specific subsets of vertices. However, existing approaches [8, 15] of fitting graph data into tabular data often treat labels as regular properties, encapsulating them within a string or list, as seen in Figure 3. These approaches overlook the inherent differences between labels and other properties, leading to inefficiencies when performing filtering operations based on label conditions, due to the need for decoding string representations and string matching.
@@ 5.1 Handling Simple Conditions
Encoding. To compress consecutive runs of 0s or 1s, we utilize the technique of run-length encoding (RLE), which represents them as a single number. This run-length format naturally transforms the binary representation of a label into an interval-based structure. Consequently, we adopt a list denoted as $P$, to define the positions of intervals. Notably, the $i$-th interval is represented by $[P[i],P[i+1])$, where $P[i]$ refers to the $i$-th element within the list. Besides, it is required to record whether the vertices in the first interval (denoted as $[P[0],P[1])$) contain the label or not (i.e., the first value). By leveraging this technique, the storage consumption of labels can be significantly reduced.
@@ 5.2 Extending to Complex Conditions
Expanding beyond the realm of simple label existence, we encounter the intricacies of dealing with complex conditions involving multiple labels. Consider a scenario where we need to find vertices with specific label combinations, such as the GQL pattern MATCH (person:Asian&Enrollee) (or in Cypher, MATCH (person:Asian:Enrollee)), which retrieves vertices labeled as Asian and Enrollee. A more complex pattern can be MATCH (person:(Asian&!Enrollee)|Student), which retrieves vertices labeled as Asian but not Enrollee, or labeled as Student. To handle such scenarios, we employ user-defined functions (UDFs) to represent complex filter conditions. Specifically, the UDF $f$ takes a vertex $v$ as input and returns a boolean value $f(v)$, indicating whether the vertex satisfies the condition or not. Formally, we define the complex condition filtering as follows.
@@
Consider $k$ interval lists $P_{0},P_{1},\dots,P_{k-1}$, where the vertices in an interval $[P_{i}[j],P_{i}[j+1])$ share the same value for the $i$-th label. Then, if no position $P_{i}[j]$ breaks the interval $[s,e)$, i.e.,
@@ 6 Evaluation
In this section, we evaluate GraphAr on a range of real-world and synthetic graphs. We begin with micro-benchmarks that assess neighbor retrieval and label filtering, followed by an exploration of end-to-end workloads. Our findings validate GraphAr as an efficient storage scheme for LPG storage and querying in data lakes.

### 21 Dec 2023 v3

# Enhancing Data Lakes with GraphAr: Efficient Graph Data Management with a Specialized Storage Scheme

@@ 1 Introduction
In sync with these trends, graph data has become increasingly importance, especially for modeling complex relationships. Leading graph databases like Neo4j [13], TigerGraph [31], and JanusGraph [10] leverage the Labeled Property Graph (LPG) model [20, 23, 21] for this purpose. The recent ISO SQL:2023 standard includes a SQL/PGQ extension that not only facilitates querying property graphs but also enables the creation of property graph views from relational tables [19]. This groundbreaking inclusion highlights the growing convergence of relational and graph data models and emphasizes the need to integrate LPGs into data lakes. Consequently, LPGs are making their way into data lakes for multiple uses, from backups and archival storage for existing graph databases to natural extensions of transactional, log, and tabular data.
Managing and analyzing LPGs in data lakes offers several significant benefits. Graph-specific queries are often more naturally articulated in languages like Cypher [35], Gremlin [55], GQL [34], or SQL/PGQ, providing an intuitive framework for conducting comprehensive analysis of entity relationships, facilitating the discovery of valuable insights. Data lakes also provide computational flexibility for running complex graph algorithms, enabling the exploration of intricate patterns. Additionally, they offer cost-effective storage solutions, allowing organizations to utilize more affordable and colder storage options without sacrificing query performance. Most notably, data lakes enable seamless querying across both graph and relational data, ushering in a holistic approach to data analytics.
@@ 2.1 Tabular Formats in Data Lakes
Tabular data is key to data lakes, aiding efficient organization, analysis, and data extraction from large sets. Columnar formats like Parquet [5] and ORC [4] are popular due to their robust features. Unlike row-based formats such as CSV, they allow faster queries by enabling selective column reading, avoiding unnecessary data. Additionally, they offer diverse and efficient compression and encoding strategies, such as delta encoding to compress the variance between consecutive values, and run-length encoding to compress repetitive values. These techniques not only reduce storage needs but also enhance processing speeds. Another advantage of Parquet and ORC is predicate pushdown, which enhances query performance by moving filters closer to the storage layer, thereby reducing subsequent data reads.
The combination of selective column reading, efficient compression, and predicate pushdown positions Parquet and ORC as the go-to choices for managing tabular data in data lakes. Previous studies have demonstrated the importance of leveraging their capabilities for optimizing relational data management [33, 40, 22]. Recent research [45, 60, 61, 62, 63, 6] has also explored enhancements to tabular formats, utilizing CPU instruction sets like BMI and SIMD. In this paper, we will focus on Parquet, but the techniques discussed can be seamlessly adapted to other columnar formats such as ORC.
@@ 2.2 Labeled Property Graphs
Labeled Property Graphs (LPGs) [20, 23, 21] excel at representing complex relationships and semantics in a natural manner. Their flexible schema allows for accommodating the diverse and evolving nature of big data within repositories, making them integral to data lakes. LPG serves as the canonical data model in many graph systems [13, 36, 31] and graph query languages [59, 7, 35, 55, 34, 30], enabling queries and analytics to uncover valuable insights and patterns. Formally, an LPG is defined as $G=(V,E,T_{V},T_{E},P,L)$, where $V$ is a set of vertices, $E$ a set of edges, $T_{V}$ and $T_{E}$ the types of vertices and edges respectively, $P$ the properties, and $L$ the labels.
@@
Furthermore, each label $l\in L$ has a unique identifier, usually a string. Each vertex type $t_{v}$ is linked with a set of candidate labels $L(t_{v})\subseteq L$, allowing each vertex of this type to be assigned zero or more labels. Labels hold significant importance in LPGs as they represent classifications and characteristics of entities, whereas properties serve as attributes to store additional information. Although edge types could technically also be labeled, in this paper, we focus solely on vertex labels, aligning with common graph query practice  [^1] Graph query languages like Cypher and Gremlin typically adhere to the convention that an edge can have only one classification, corresponding to the edge type in LPG model. Nevertheless, the strategies for vertex labels discussed in this paper can be seamlessly extended to support edge labels..
@@ 2.3 Querying Labeled Property Graphs
These query-side inefficiencies highlight the limitations of using tabular formats for graphs. The reliance on suboptimal workarounds, such as multiple joins for topology expansion and makeshift encoding schemes for label filtering, compromises query performance and complicates the query process. This paves the way for the challenges we seek to address.
@@ 2.4 Key Challenges Addressed by GraphAr
The development of GraphAr is motivated by the specific limitations of existing tabular formats for both representing LPGs and supporting efficient graph queries.
@@
Each of these challenges represents a gap in the capabilities of current tabular formats for graph data. They serve as the focus areas for the technical contributions of GraphAr, with each corresponding to a dedicated section in this paper.
(a) Metadata.
(b) Vertex table of Person (and Disease).
(c) Edge table of Person-Diagnosed-Disease.
@@ 3.1 Goals and Non-Goals
Non-Goals. GraphAr does not intend to replace existing data lake formats like Parquet and ORC, but to maximize their benefits and offer additional features for LPGs.
@@ 3.2 Data Organization and Layout
Property columns (pid and age) are named after their respective properties and hold the corresponding values with specified datatypes. In terms of labels, a set of candidate labels is defined for each vertex type. Then a vertex can have an arbitrary number of labels from the corresponding set. For example, the vertex type Person may have labels to represent ethnicity. For efficient storage and filtering of labels, GraphAr uses a binary representation to maintain each label in an individual column named with angle brackets, e.g., <Asian> and <Enrollee>. Additionally, advanced encoding/decoding techniques are applied, which will be discussed in Section 5.
Edge table. Edges are also organized and stored in Parquet files, similar to vertices. Figure 3(c) showcases the layout of the edge table for type Person-Diagnosed-Disease, where Person and Disease represent the source and destination vertex types, while Diagnosed signifies the classification of the relationships. Each edge is associated with the internal IDs of its source and destination vertices, stored in columns named <src> and <dst>. Edge properties, and optional partitioning, are handled in the same way as vertices.
@@ 3.3 Incorporation with Data Lakes
The design of GraphAr makes it especially well-suited for integration with data lakes, largely due to its reliance on widely adopted standards like Parquet and YAML.
Data transformation and construction. The GraphAr format is essentially a specialized layout of Parquet files accompanied by a YAML metadata file. This enables the use of existing data processing frameworks like Apache Spark, Acero, and Hadoop, which can access various graph systems like Neo4j, TigerGraph and Nebula, or other types of database systems through their respective connectors. These frameworks can also ingest a multitude of data formats including logs, relational tables, JSON, and more. Such flexibility provides users with the ability to construct, transform, and store LPGs in data lakes from a wide array of data sources. To simplify the process of generating files in GraphAr, we have also provided a Spark library specifically designed for this purpose.
@@ 4.1 Workflow and Challenges
Given a column in vertex table that includes $m$ pages, the PAC $\mathcal{C}=[C_{0},\ldots,C_{m-1}]$ is a list of up to $m$ collections. Each $C_{i}$ stores a set of internal IDs in the corresponding page. Non-empty collections in $\mathcal{C}$ are retained, while empty ones are omitted.
@@ 4.2 Delta Encoding
Implementations. We utilize Parquet’s built-in support for delta encoding [43], which is implemented based on miniblocks. Each miniblock (with a size of $32$ values) is binary packed using its own bit width, which should be a power of 2 for data alignment purposes. This design allows us to adapt to changes in the data distribution, as the bit width of each miniblock is dynamically adjusted to minimize storage consumption. According to our evaluation across various real-world graphs, as detailed in Section 6.2, the delta encoding technique can reduce the expected loaded data volume (which can be measured by storage consumption) by $58.1\%$ to $81.0\%$ compared to without delta encoding. As a result, it brings an individual speedup of up to 3.6×3.6$\times$3.6 × for neighbor retrieval.
@@ 4.3 BMI-based Decoding
Another unsolved issue is the choice of implementing collection $C$ in PAC. A pioneering solution proposed in [45] underscores the transformation of indices into a bitmap representation to enable selection pushdown in columnar storages. To adopt this approach to efficiently retrieve the properties of neighbors, which constitutes the second task of neighbor retrieval, we adopt a bitmap representation $B$ for each non-empty collection $C$ in PAC, where $B[i]=1$ indicates the existence of $i$-th element. Figure 5 illustrates a PAC and its bitmap representation, where only $C_{1}$ is non-empty, and $B_{1}$ is the bitmap representation of $C_{1}$. The bitmap representation can be used to facilitate the selection pushdown of vertex properties or labels, e.g., fetching the properties of name in the target vertex table Disease, for the neighbors of a person vertex.
@@
Intuition. To ensure clarity, we initially consider a simplified scenario where each delta value is compressed to a 4-bit size. Conventionally, a two-step approach is used to decode the delta-encoded data, in which the current encoded value is added to the previously decoded ID to obtain the current ID, and then the bitmap is updated bit by bit based on the decoded IDs. However, our analysis reveals that this two-step process is redundant. By leveraging the bit-shifting encoding $1<<(d-1)$ for each delta value $d$, we can generate the bitmap by concatenating the bit-shifting encodings: $1<<(d_{n-1}-1)||\dots||1<<(d_{1}-1)||1<<(d_{0}-1)$, where $d_{0},d_{1},\dots,d_{n-1}$ represent $n$ delta values, and ||||| | represents the concatenation operator. This principle is visually depicted in Figure 5(a).
@@
The subsequent challenge is to generate the required mask, achieved by deriving the $i$-th mask $m_{i}$ from the $i$-th bit-shifting encoding $s_{i}$ using the equation $m_{i}=(s_{i}<<1)-1$. The beauty of this operation is its potential for parallel execution, facilitated by direct manipulation of the mask sequence residing within a 64-bit register. Specifically, $M$ can be acquired via the following steps: 1) a bitwise (in our example, 16 bits) shift of each $s_{i}$ to the left by 1 bit, parallelized through SIMD instructions like _mm_slli_epi16; and 2) a bitwise subtraction of 1 from the result of the previous step, which can be accelerated via instructions like _mm_sub_epi16. All these SIMD instructions utilized in our implementations are widely available in modern CPUs, included in SSE2 (which we use) and more recent sets such as AVX2 and AVX-512.
@@ 5 Optimized Label Filtering
Labels serve as a representation of the classification or characteristics of vertices in a graph. Filtering vertices by labels is a fundamental syntax in graph query languages, as it allows querying specific subsets of vertices. However, existing approaches [8, 15] of fitting graphs into tabular data often treat labels as regular properties, encapsulating them within a string or list, as seen in Figure 3. These approaches overlook the inherent differences between labels and properties, leading to inefficiencies when performing label filtering, due to the need for decoding string representations and string matching.
@@ 5.1 Handling Simple Conditions
Encoding. To compress consecutive runs of 0s or 1s, we utilize the technique of run-length encoding (RLE), which represents them as a single number. This run-length format naturally transforms the binary representation of a label into an interval-based structure. We then adopt a list $P$ to define the positions of intervals. The $i$-th interval is represented by $[P[i],P[i+1])$, where $P[i]$ refers to the $i$-th element within $P$. Besides, it is required to record whether the vertices of the first interval $[P[0],P[1])$ contain the label or not, i.e., the first value. By leveraging this technique, the storage consumption of labels can be significantly reduced.
@@ 5.2 Extending to Complex Conditions
Expanding beyond the realm of simple label existence, we encounter the intricacies of dealing with complex conditions involving multiple labels. Consider a scenario where we need to find vertices with specific label combinations, such as the GQL pattern MATCH (person:Asian&Enrollee) (or in Cypher, MATCH (person:Asian:Enrollee)), which retrieves vertices labeled as Asian and Enrollee. A more complex pattern can be MATCH (person:(Asian&!Enrollee)|Student), which retrieves vertices labeled as Asian but not Enrollee, or labeled as Student. To handle such scenarios, we employ user-defined functions (UDFs) to represent complex filter conditions. The UDF $f$ takes a vertex $v$ as input and returns a boolean value $f(v)$, indicating whether the vertex satisfies the condition or not. Formally, we define the complex condition filtering as follows.
@@
Consider $k$ interval lists $P_{0},P_{1},\dots,P_{k-1}$, where the vertices in $[P_{i}[j],P_{i}[j+1])$ share the same value for the $i$-th label. If an iterval $[s,e)$ is not broken by any position, i.e.,
@@ 6 Evaluation
In this section, we evaluate GraphAr on a range of real-world and synthetic graphs, through micro-benchmarks focusing on neighbor retrieval and label filtering, as well as end-to-end workloads. Our findings validate GraphAr as an efficient storage scheme for LPG storage and querying in data lakes.